OpenStudy (meilendurcer):

Evaluate. A. -5 B. 3 C. 15 D. 25

OpenStudy (meilendurcer):

\[\sqrt{49}+4^{2}-3-\sqrt{25}\]

OpenStudy (anonymous):

\[\sqrt{49} + 4^2 - 3 - \sqrt{25}= 7 + 16 - 3 - 5\]